OpenStudy (anonymous):

Two friends went to the mall together and the held each others bags. One friend found change in the bottom of the bag. They both claimed the change. The friend that found the change said that if the other friend can guess the exact coins that she found she can keep the change if not the one who found it gets to keep it. Here are the clues, there are twenty coins all together, and there are only pennies and nickels. The total value of the coins is 48 cent. How many pennies and how many nickels did the friend find?

OpenStudy (anonymous):

5x+y=48 Thats the equation for total value x+y=20 Thats the equation for the amount of coins Solve the system of equations x=nickels y=pennies

OpenStudy (anonymous):

isnt 5x+y=48 a function?

OpenStudy (anonymous):

you could make it a function, a function has to be y= but thats not the point, ya gotta solve the system to get x and y

OpenStudy (anonymous):

7 nickles and 13 pennies

OpenStudy (anonymous):

I have to write a report on how I got my answer, so would I put the equations you put?

OpenStudy (anonymous):

yeah